technician compares repair costs for two types of microwave ovens (type I and type II). He believes that the repair cost for type I ovens is greater than the repair cost for type II ovens. A sample of 57 type I ovens has a mean repair cost of $82.19 , with a standard deviation of $11.01 . A sample of 52 type II ovens has a mean repair cost of $80.18 , with a standard deviation of $22.47 . Conduct a hypothesis test of the technician's claim at the 0.05 level of significance. Let μ1 be the true mean repair cost for type I ovens and μ2 be the true mean repair cost for type II ovens.

Step 1 of 4:

State the null and alternative hypotheses for the test.

Step 2 of 4:

Compute the value of the test statistic. Round your answer to two decimal places.

Step 3 of 4:

Determine the decision rule for rejecting the null hypothesis H0 . Round the numerical portion of your answer to three decimal places.

Step 4 of 4:

Make the decision for the hypothesis test.
Answer #1

Ans:

1)

2)Test statistic:

t=(82.19-80.18)/sqrt((11.01^2/57)+(22.47^2/52))

t=0.584

3)df=52-1=51

critical t value=1.675

4)Fail to reject the null hypothesis.